HUBBLE SPACE TELESCOPE DAILY REPORT #1416 PERIOD COVERED: 0000Z (UTC) 06/20/95 - 0000Z (UTC) 06/21/95 Daily Status Report as of 172/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.2 Completed FOS/BL 5539 (Cycle 4 FOS Photometric Monitor) The Faint Object Spectrograph (Blue Side) used standard star BD+28D4211 to calibrate the absolute sensitivity of its detectors. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 6251 (Medium Deep Survey - Crafting Test) The WF/PC-2 was used in this continuation of the medium deep survey. The proposal is targeting rare objects (lenses, peculiar galaxies, faint stars) and sources related to the faint blue galaxy problem. This observation was completed with no reported problems. 1.4 Completed WF/PC-2 5778 (WF/PC-2 Cycle 4 Lyman Alpha Throughput Verification Test) The WF/PC-2 was used to make ultraviolet observations of standard star BD+75D325 through filter F122M as part of a Lyman Alpha throughput verification test. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed GHRS/1 5542 (GHRS Side One Wavelength (1st Order Gratings): Cycle 4) The High Resolution Spectrograph (Side One) used the spectral calibration lamp (Side Two) to calibrate the Side One characteristics. Notice that the Side Two lamp was used to calibrate Side One. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Notes: There were three instances yesterday afternoon (of approximately 47 minutes, 36 minutes and 26 minutes) of real time engineering data loss due to STGT outage. TTR 18425 documents these events. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-